Do the following work for a set of quantum numbers
Answer #1

a) does not work

because l values start with n-1

if n = 2

l must be start with 1

but in set a n = l it is not acceotable

b) yes it works

n = 1

l = 1 - 1 = 0

ml = 0

ms = -1/2